### v1

# A note on strong similarity and the Connes embedding problem

Abstract
We show that there exists a completely bounded (c.b. in short) homomorphism $u$ from a $C^{*}$-algebra $C$ with the lifting property (in short LP) into a QWEP von Neumann algebra NN that is not strongly similar to a *-homomorphism, i.e. the similarities that “orthogonalize” $u$ (which exist since $u$ is c.b.) cannot belong to the von Neumann algebra NN. Moreover, the map $u$ does not admit any c.b. lifting up into the WEP $C^{*}$-algebra of which NN is a quotient. We can take $C=C^{*}(\mathbb{F})$ (full $C^{*}$-algebra) where $\mathbb{F}$ is any nonabelian free group and $N=B(H)\bar{\otimes}M$ where $M$ is the von Neumann algebra generated by the reduced $C^{*}$-algebra of $\mathbb{F}$.
In this note (which is not intended for publication) we produce a c.b. homomorphism $u$ from a $C^{*}$-algebra $C$ with LP (which we could choose to be $C^{*}(\mathbb{F}_{n})$ if we wish using free Haar unitaries in place of $(x_{j})$) into a QWEP von Neumann algebra NN which is NOT strongly similar to a *-homomorphism, here by strongly similar we mean that the similarity can be chosen in the von Neumann algebra NN.
This means that a certain statement from [6] crucially used in [1] is not valid, as well as the application proposed in [1] to the Connes embedding problem.
If $N=B/I$ with $B$ WEP (II closed two sided ideal in $B$) the map $u$ does not admit any c.b. lifting. Indeed the paper [1] shows that if it did then $u$ would be strongly similar to a *-homomorphism.
Remark 0.1.
By known results (see [2]) there are c.b. maps $u:C\to B/I$ that do not admit any c.b. lifting. Apparently the existence of a c.b. homomorphism is new.
Let $\ell_{j}$ (resp. $r_{j}$) denote the left (resp. right) creators on the full Fock space $H$ so that
$x_{j}=\ell_{j}+\ell_{j}^{*}\quad(1\leq j\leq n)$
are free semicirculars generating the von Neumann algebra $M\subset B(H)$ and their commutant $M^{\prime}$ is generated by
$y_{j}=r_{j}+r_{j}^{*}.$
Let $C=\ast_{j}C^{*}(x_{j})$ (full free product). Let $\sigma:C\to M$ be the natural representation, so that $M=\sigma(C)^{\prime\prime}$, and let $\pi:C\to M\bar{\otimes}B(H)$ be defined by $\pi(x)=\sigma(x)\otimes 1$.
Let $N=B(H)\bar{\otimes}M$. We will produce a c.b. derivation $\delta:C\to N$ relative to $\pi$ with $\|\delta\|_{cb}\leq 2$ but such that any $T$ in NN such that $\delta(x)=[T,\pi(x)]$ satisfies $\|T\|\geq\sqrt{n}/4$.
Then we will show that the associated homomorphism $u:C\to M_{2}(N)$ defined by
$C\ni x\mapsto\left(\begin{matrix}x\otimes 1&\ \delta(x)\\ 0&\ x\otimes 1\end{matrix}\right)$
is also c.b. but if $n=\infty$ it is not strongly similar to a *-homomorphism. An estimate will also be given for large finite $n$’s.
The construction we use is a particular case of the one described in [4].
Let
$S=\sum-r_{j}\otimes\ell_{j}^{*}+r_{j}^{*}\otimes\ell_{j}\in B(H\otimes H)$
Clearly $\|\sum\ell_{j}\ell_{j}^{*}\|=\|\sum r_{j}r_{j}^{*}\|\leq 1$ and hence
$\|S\|\leq 2.$
Let $C=\ast_{j}C^{*}(x_{j})$ (full free product) We denote by $\pi:C\to M\otimes 1$ the representation from the full free product to $M$ taking $x_{j}$ to $x_{j}\otimes 1$ .
Note that $C$ has the LP.
Let $\delta:C\to B(H\otimes H)$ be defined by
$\delta(x)=[x\otimes 1,S].$
Then $\delta$ is inner with
(0.1) $\|\delta\|_{cb}\leq 2.$
Since $x_{i},y_{j}]=0$ we have
$\delta(x)=[x\otimes 1,\sum r_{j}^{*}\otimes\ell_{j}^{*}+r_{j}^{*}\otimes\ell_{j}]=[x\otimes 1,\sum r_{j}^{*}\otimes x_{j}]=\sum[x,r_{j}^{*}]\otimes x_{j}=-\sum[x,r_{j}]\otimes x_{j}.$
Clearly the von Neumann algebra generated by $\pi(C)$ and $\delta(C)$ is included in $B(H)\bar{\otimes}\{x_{j}\}^{\prime\prime}=B(H)\bar{\otimes}M$.
Let $T\in B(H)\bar{\otimes}M$ be such that for all $x\in C$
$\delta(x)=[x\otimes 1,T].$
Since we have a projection $P$ on $M$ with range the span of $(x_{j})$ with $\|P\|_{cb}\leq 2$ we must have
$\delta(x)=(Id\otimes P)\delta(x)=[x\otimes 1,T_{1}]$
with
$T_{1}=(Id\otimes P)(T).$
It follows that $T_{1}=\sum z_{j}\otimes x_{j}$ and
(0.2) $\max\{\|\sum z_{j}z_{j}^{*}\|^{1/2},\|\sum z_{j}^{*}z_{j}\|^{1/2}\}\leq\|T_{1}\|\leq 2\|T\|.$
Moreover since
$\delta(x)=\sum[x,r_{j}^{*}]\otimes x_{j}=-\sum[x,r_{j}]\otimes x_{j}=\sum[x,z_{j}]\otimes x_{j}$
we must have
[…]
$y_{j}=(r_{j}+z_{j})+(r_{j}^{*}-z_{j})$
and hence
$\sqrt{n}=(\sum\|y_{j}\Omega\|^{2})^{1/2}\leq(\sum\|(r_{j}+z_{j})\Omega\|^{2})^{1/2}+(\sum\|(r_{j}^{*}-z_{j})\Omega\|^{2})^{1/2}$
but by the trace identity in $M^{\prime}$
$(\sum\|(r_{j}+z_{j})\Omega\|^{2})^{1/2}=(\sum\|(r_{j}+z_{j})^{*}\Omega\|^{2})^{1/2}=(\sum\|z_{j}^{*}\Omega\|^{2})^{1/2}$
and clearly $(\sum\|(r_{j}^{*}-z_{j})\Omega\|^{2})^{1/2}=(\sum\|z_{j}\Omega\|^{2})^{1/2}$ so that we obtain by (0.2)
(0.3) $\sqrt{n}\leq\|\sum z_{j}z_{j}^{*}\|^{1/2}+\|\sum z_{j}^{*}z_{j}\|^{1/2}\leq 2\|T_{1}\|\leq 4\|T\|.$
Now it remains to check that the von Neumann algebra generated NN by $\pi(C)$ (or $M\otimes 1$) and $\delta(C)$ is QWEP.
We will show that $N=B(H)\bar{\otimes}M$, which is clearly QWEP since $M$ being QWEP is exactly the original observation of Connes in the 1976 paper where he raised what is now known as the Connes embedding problem.
Recall
$\delta(x)=\sum[x,r_{j}^{*}]\otimes x_{j}=-\sum[x,r_{j}]\otimes x_{j}$
Observe that $l_{k},r_{j}]=[l_{k}^{*},r_{j}^{*}]=0$ for all $k,j$ and also $l_{k},r_{j}^{*}]=0$ if $k\not=j$ and $l_{k},r_{k}^{*}]=P_{\Omega}$. Thus
$\delta(x_{k})=P_{\Omega}\otimes x_{k}$
which implies that NN contains all elements of the form $aP_{\Omega}b\otimes x_{k}$ ($a,b\in M$), i.e. all elements in $F(H)\otimes x_{k}$ where $F(H)$ denotes the space of finite rank operators on $H$. It follows that NN contains $F(H)\otimes M$ and hence $N\supset B(H)\bar{\otimes}M$. Since we already noted that the converse is obvious, this proves as announced $N=B(H)\bar{\otimes}M$.
The arguments in [1] give the following:
Proposition 0.2.
Consider $u:C\to N$ a unital cb homomorphism with values in N assumed to be a QWEP von Neumann algebra in $B(H)$. Let $v:N^{\prime}\to B(H)$ be the inclusion map. Then $u$ is strongly similar to a *-homomorphism iff the map $u.v:C\otimes N^{\prime}\to B(H)$ (defined by $u.v(a\otimes b)=u(a)v(b)$ extends to a c.b. map $w$ from $C\otimes_{\max}N^{\prime}$ to $B(H)$. Moreover
$\|w:C\otimes_{\max}N^{\prime}\to B(H)\|_{cb}=\inf\{\|T\|\|T^{-1}\|\}$
where the (attained) inf runs over all invertible $T\in N$ such that $Tu(.)T^{-1}$ is a *-homomorphism.
Proof.
Clearly $w$ is still a homomorphism, and hence if it is cb it is similar to a *-homomorphism whence the existence of an invertible $T$ on $H$ (with $\|T\|\|T^{-1}\|=\|w\|_{cb}$) such that $x\mapsto Tw(x)T^{-1}$ is a *-homomorphism. But restricting this to $1\otimes N^{\prime}$ we find that $N^{\prime}\ni y\mapsto TyT^{-1}$ is a *-homomorphism which implies by a simple calculation ($TyT^{-1}=(Ty^{*}T^{-1})^{*}$ and hence $T^{*}Ty=yT^{*}T$) that $|T|\in N^{\prime\prime}=N$. Now $T=U|T|$ with $U$ unitary and the map $C\ni x\mapsto U(|T|u(x)|T|^{-1})U^{-1}$ is a *-homomorphism. Since $U$ is unitary $C\ni x\mapsto(|T|u(x)|T|^{-1})$ is a *-homomorphism, proving the strong similarity of $u$.
Conversely, if $T\in N$ is such that $C\ni x\mapsto(Tu(x)T^{-1})$ is a *-homomorphism denoted by $\rho:C\to N$, then obviously $\rho.v$ is a (completely contractive) *-homomorphism from $C\otimes_{\max}N^{\prime}$ to $B(H)$. Then $a\otimes b\mapsto T^{-1}\rho.v(a\otimes b)T$ has cb-norm at most $\|T\|\|T^{-1}\|$ on $C\otimes_{\max}N^{\prime}$ and since $T$ commutes with $v(b)$, we have $T^{-1}\rho.v(a\otimes b)T=u.v(a\otimes b)$. Thus we conclude that $\|u.v:C\otimes_{\max}N^{\prime}\to B(H)\|_{cb}\leq\|T\|\|T^{-1}\|$. The converse inequality was proved in the first part. ∎
Let $\delta$ be as before. Then
$C\ni x\mapsto u(x)=\left(\begin{matrix}x\otimes 1&\ \delta(x)\\ 0&\ x\otimes 1\end{matrix}\right)$
is a c.b. homomorphism $u$ into $M_{2}(N)$.
We have $\|u\|_{cb}\leq 1+\|\delta\|_{cb}\leq 3$ by (0.1) (and replacing $\delta$ by $(\varepsilon/2)\delta$ we could get $\|u\|_{cb}\leq 1+\varepsilon$ if desired).
We know that $M_{2}(N)$ is QWEP, so we have $M_{2}(N)=B/I$ with $B$ WEP.
Lemma 0.3.
Let $S$ be an invertible in $M_{2}(N)$ such that $x\mapsto Su(x)S^{-1}$ is a *-homomorphism. Then there is $T$ in NN with
$\|T\|\leq(\|S\|\|S^{-1}\|)^{2}$
such that $\delta(x)=[x,T]$ for any $x\in C$.
Proof.
See [3, p. 80] where this is proved following an argument of Paulsen. ∎
The following, which actually is an interesting fact, shows that [6, Lemma 2.3] is not correct :
Corollary 0.4.
The above c.b. homomorphism $u:C\to M_{2}(N)$, which has
$\|u\|_{cb}\leq 3$
is such that any c.b. lifting $\hat{u}:C\to B$ satisfies
$\sqrt{n}/4\leq(\|\hat{u}\|_{cb})^{2}.$
Thus if we work with $n=\infty$ we obtain a $u$ without any c.b. lifting.
Proof.
Indeed, it would follow (here $\mathscr{C}=C^{*}(\mathbb{F}_{\infty})$)
$\|Id_{\mathscr{C}}\otimes\hat{u}:\mathscr{C}\otimes_{\min}C\to\mathscr{C}\otimes_{\min}B\|\leq\|\hat{u}\|_{cb}$
and since $B$ WEP implies $\mathscr{C}\otimes_{\min}B=\mathscr{C}\otimes_{\max}B$, we would have
$\|Id_{\mathscr{C}}\otimes\hat{u}:\mathscr{C}\otimes_{\min}C\to\mathscr{C}\otimes_{\max}B\|\leq\|\hat{u}\|_{cb}$
and a fortiori
$\|Id_{\mathscr{C}}\otimes\hat{u}:\mathscr{C}\otimes_{\max}C\to\mathscr{C}\otimes_{\max}B\|\leq\|\hat{u}\|_{cb}.$
Since $u=q\hat{u}$, we also have (recall $M_{2}(N)=B/I$)
$\|Id_{\mathscr{C}}\otimes u:\mathscr{C}\otimes_{\max}C\to\mathscr{C}\otimes_{\max}M_{2}(N)\|\leq\|\hat{u}\|_{cb}.$
With the notation in [5] this means
$\|u\|_{mb}\leq\|\hat{u}\|_{cb},$
which implies
$\|Id_{M_{2}(N)^{\prime}}\otimes u:{M_{2}(N)^{\prime}}\otimes_{\max}C\to{M_{2}(N)^{\prime}}\otimes_{\max}M_{2}(N)\|_{cb}\leq\|\hat{u}\|_{cb}.$
So (after flipping the factors) denoting as before by $V:{M_{2}(N)^{\prime}}\to M_{2}(B(H))$ the inclusion map, we obtain a fortiori
$\|u.V\|_{cb}\leq\|\hat{u}\|_{cb}.$
By the proposition and the lemma above there is an invertible $T$ in NN with
$\|T\|\|T^{-1}\|\leq(\|\hat{u}\|_{cb})^{2}$
such that $\delta(x)=[x,T]$ for any $x\in C$. By (0.3) this implies
$\sqrt{n}/4\leq(\|\hat{u}\|_{cb})^{2}.$
Remark 0.5 (Free group variant).
Let $(g_{j})$ be the free generators of a free group $G$. Let $\lambda$ and $\rho$ denote the left and right regular representations on $G$. It is well known (see [4]) that $\rho(g_{j})$ can be written as $a_{j}+b_{j}$ with $\sum a_{j}^{*}a_{j}\leq 1$ and $\sum b_{j}b_{j}^{*}\leq 1$ (and of course similarly for $\lambda(g_{j})$). The preceding construction runs just the same with $C=C^{*}(G)$ if one takes for $M$ the von Neumann algebra generated by $\lambda$ and one sets $\delta(x)=\sum[x,a_{j}]\otimes\lambda(g_{j})$ for $x\in M$. In other words we may replace $(x_{j})$ (free gaussians) by $\lambda(g_{j})$ (free Haar unitaries) with no essential change.
Appendix A close inspection of [6] shows that the error lies in an incorrect identification. Using the notation of [6] the identity
(0.4) $TBT^{-1}/J=S(B/J)S^{-1}$
is incorrect.
Indeed, we have
$B=\{pb+(1-p)b\mid b\in B\}\subset B^{**}\subset B(H)$
with
$H=pH\oplus(1-p)H$
and
$TBT^{-1}\subset B(H),$
which can be described as follows:
$TBT^{-1}=\{S(pb)S^{-1}+(1-p)b\mid b\in B\}=\{y_{1}+y_{2}\mid S^{-1}(y_{1})S+y_{2}\in B,\ y_{1}\in pB^{**},y_{2}\in(1-p)B^{**}\}.$
Indeed, assuming $y_{1}\in pB^{**},y_{2}\in(1-p)B^{**}$, if $b=S^{-1}(y_{1})S+y_{2}$ then $pb=S^{-1}(y_{1})S$ and $(1-p)b=y_{2}$ so that $y_{1}+y_{2}=S(pb)S^{-1}+(1-p)b$. This proves $\supset$”. The reverse is clear since $pB\subset pB^{**}$ and $(1-p)B\subset(1-p)B^{**}$.
Note
$J=(1-p)J=\{(1-p)j\mid j\in J\}.$
Since $B/J\simeq pB$ the identity (0.4) must be interpreted as $TBT^{-1}/J=S(pB)S^{-1}$. This implies
$TBT^{-1}\subset S(pB)S^{-1}+J,$
and hence after the action of $T^{-1}(.)T$
$B\subset pB+J$
which implies
$(1-p)B\subset(1-p)J=J.$
However, in general we have $(1-p)B\not=J$. To verify that, assume (1p)B=J(1-p)B=J. Let $f:B\to J$ be defined by f(x)=(1p)x.f(x)=(1-p)x. Then $f$ is a *-homomorphism and a linear projection from $B$ to (1p)B=J(1-p)B=J. In otherwords we have a splitting of $B$ as $B/J\oplus J$, which of course fails in general.

### v2

@@
 Proposition 0.2.
-Consider $u:C\to N$ a unital cb homomorphism with values in N assumed to be a QWEP von Neumann algebra in $B(H)$.
+Consider $u:C\to N$ a unital cb homomorphism with values in a von Neumann algebra N in $B(H)$.
 Let $v:N^{\prime}\to B(H)$ be the inclusion map.
@@
 Proof.
-Clearly $w$ is still a homomorphism, and hence if it is cb it is similar to a *-homomorphism whence the existence of an invertible $T$ on $H$ (with $\|T\|\|T^{-1}\|=\|w\|_{cb}$) such that $x\mapsto Tw(x)T^{-1}$ is a *-homomorphism.
+Assume $w$ c.b.
+Clearly $w$ is still a homomorphism, and hence if it is c.b.
+it must be similar to a *-homomorphism whence the existence of an invertible $T$ on $H$ (with $\|T\|\|T^{-1}\|=\|w\|_{cb}$) such that $x\mapsto Tw(x)T^{-1}$ is a *-homomorphism.
 But restricting this to $1\otimes N^{\prime}$ we find that $N^{\prime}\ni y\mapsto TyT^{-1}$ is a *-homomorphism which implies by a simple calculation ($TyT^{-1}=(Ty^{*}T^{-1})^{*}$ and hence $T^{*}Ty=yT^{*}T$) that $|T|\in N^{\prime\prime}=N$.